OpenStudy (anonymous):

If f(x) = -5x + 1, find f(-3).

OpenStudy (anonymous):

16

OpenStudy (anonymous):

yes its easy question

OpenStudy (anonymous):

thanksss :D ! can you show how you got it though?

OpenStudy (anonymous):

f(x)=-5x+1 So f(-3)=-5(-3)+1 =15+1 =16

OpenStudy (anonymous):

f(-3)=-5*-3+1=16